OpenStudy (abbycross167):

Can someone please help me fix this math question? Please? I will fan and give a medal!! Please help me to simplify this rational expression - (4x^2-6x-18) (x^2+1) ------------ * ---------- (x+1) (x^2-6x+9)

OpenStudy (abbycross167):

These are the Steps I did- (4x^2-6x-18)(x^2+1) ---------------------- (x+1)(x^2-6x+9) (4x^2-6x-18) = 2(2x^2-3x-9)= 2(2x^2-6x+3x-9)= 2(2x^2-6x)+(3x-9)= 2(2x^2(x-3)+3(x-3) 2(x-3)(2x+3)= 2x^3 ------------- (x+1)(x^2-6x+9)

OpenStudy (abbycross167):

This is what my teacher said was wrong - Make your steps look like this example - (8y) (7x) ------ * ------ (14x^2) (8z^3) Steps-- 1 1 1 1 1 2*2*2*y * 7 * z' ---------------------- 2*7*x*x 2*2*2*z*z*z 1 1 1 1 1 1*1*1*y*1*1 ------------------- 2*1*1*x*1*1*1*z*z*z y ------ 2xz^3

OpenStudy (retireed):